# Towards deep physics-informed Kolmogorov–Arnold networks

### v1

@@ 5.5 Advection Equation
Table 13: Results for different RGA KAN $(\alpha,\beta)$ initializations on the advection equation. Reported values are mean $\pm$ standard error over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
The advection equation (see D for details), especially at high transport velocities (the constant multiplying the spatial derivative of the solution field in Eq. (96)), is a challenging benchmark that typically requires specialized training strategies (e.g., learnable spatial periodical embeddings as in [23]) to obtain accurate solutions. In this work, however, we intentionally refrain from introducing any problem-specific modifications to maintain a unified training pipeline across all PDEs. To this end, we set the advection velocity to $c=20$, which remains a nontrivial setting. Table 13 shows that initializing RGA KANs with $\alpha=1$ and $\beta=1$ yields the best performance, with both the lowest mean error and the smallest variability across seeds. Among these runs, the best-performing model achieves a final relative $L^{2}$ error of $1.81\times 10^{-4}$, and its prediction is depicted in Figure 13 alongside the reference solution and absolute error.
@@
Table 14: Performance comparison on the advection equation across different architectures. Reported values are mean $\pm$ standard error over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 13. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LR Annealing Relative $L^{2}$ Error
@@
Table 15: Ablation study on adaptive training components for the advection equation using the best $(\alpha,\beta)$ initialization from Table 13. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values are mean $\pm$ standard error over three seeds.
The ablation results in Table 15 highlight the importance of adaptive training for this PDE. Training with only RBA leads to divergence, as does removing both RBA and causal training (indicated by the dash in the table), or RBA and learning-rate annealing. Notably, RAD is the only adaptive strategy whose removal alongside RBA does not cause divergence, although performance is still noticeably degraded. This stands in contrast to previous benchmarks, where RAD often had the most pronounced effect, emphasizing that the relative contribution of each adaptive technique is highly problem dependent and should be evaluated individually for each PDE, as done in this study.
@@ 5.6 Helmholtz Equation
Table 16: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Helmholtz equation. Reported values are mean $\pm$ standard error over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We next consider the Helmholtz equation (see D for details), defined solely on a spatial domain and thus involving no initial condition. Consequently, this benchmark excludes both learning-rate annealing (as the loss contains only one PDE residual term) and causal training (due to the absence of temporal dependencies). Only RBA and RAD are employed as adaptive training strategies. Table 16 shows that initializing RGA KANs with $\alpha=1$ and $\beta=1$ yields the best performance, achieving both the lowest mean relative $L^{2}$ error and the smallest variance across seeds. Notably, this configuration also attains the smallest final loss of order $\mathcal{O}(10^{-4})$, which is particularly meaningful for this benchmark: the Helmholtz equation is well known to exhibit a high and challenging loss landscape at initialization [35]. The best-performing model for this configuration achieves a relative $L^{2}$ error of $1.76\times 10^{-5}$, and its prediction is shown in Figure 14.
@@
Table 17: Performance comparison on the Helmholtz equation across different architectures. Reported values are mean $\pm$ standard error over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 16. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
@@
Table 18: Ablation study on adaptive training components for the Helmholtz equation using the best $(\alpha,\beta)$ initialization from Table 16. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values are mean $\pm$ standard error over three seeds.
@@ 5.7 Poisson Equation
Table 19: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Poisson equation for $\omega\in\{1,2,4\}$. Reported values are mean $\pm$ standard error over three seeds. The best performing configuration in terms of relative $L^{2}$ error per column is indicated in bold.
The final benchmark is the Poisson equation (see D for details), which – similarly to the Helmholtz equation – is defined on a purely spatial domain and therefore involves no initial condition. As a result, only RBA and RAD are employed as adaptive training strategies. The source term is chosen such that the analytical solution is $u(x,y)=\sin(\pi\omega x)\sin(\pi\omega y)$, allowing us to systematically investigate how performance degrades as the frequency parameter $\omega$ increases. We consider $\omega\in\{1,2,4\}$ and Table 19 presents the results of the $(\alpha,\beta)$ initialization study for each $\omega$. Evidently, different initialization configurations are optimal for different frequencies, and performance deteriorates with increasing $\omega$. Moreover, the standard error of the mean for the optimal configuration grows alongside $\omega$: for $\omega=1$, it is roughly an order of magnitude lower than the mean error, whereas for $\omega=4$ it exceeds half of it, indicating increased sensitivity to initialization and optimization. For each $\omega$ value, we select the best-performing seed of the optimal configuration, achieving relative $L^{2}$ errors of $7.33\times 10^{-7}$, $2.74\times 10^{-5}$, and $3.34\times 10^{-3}$ for $\omega=1,2$, and $4$, respectively. Figure 15 shows the reference solution, prediction and absolute error for these runs.
@@
Architecture Parameters Time / Iteration (ms)
@@
Table 21: Performance comparison on the Poisson equation across different architectures. Reported values are mean $\pm$ standard error over three random seeds for $\omega\in\{1,2,4\}$. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 19. The best performing architecture in terms of relative $L^{2}$ error per column is indicated in bold.
@@
Table 22: Ablation study on adaptive training components for the Poisson equation using the best $(\alpha,\beta)$ initialization from Table 19. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values are mean $\pm$ standard error over three seeds for $\omega\in\{1,2,4\}$.
@@ 6 Conclusion and Outlook
In this work, we studied the deep training of Chebyshev-based KANs with the goal of improving their stability and accuracy in PDE benchmarks under a uniform training setup. We began by examining their initialization properties, since initialization has historically been the starting point for enabling depth scaling in neural architectures. To this end, we proposed a Glorot-like initialization scheme that is basis-agnostic and, as demonstrated by its application to sine-based KANs within the RGA KAN architecture, not tied to a specific basis family. Preliminary results on function fitting and PDE tasks showed that this initialization alone significantly improved training outcomes compared to the default initialization, in some cases by several orders of magnitude. For certain benchmarks such as Burgers’ equation, the proposed initialization was sufficient to train deeper models successfully, whereas for others, such as Allen–Cahn, it fell short. This observation made it clear that initialization alone was not enough to fully address the depth-scaling issue.
@@
Motivated by this, and inspired by PirateNets, we analyzed the properties of Chebyshev KANs at initialization and observed strong parallels to standard MLPs. This insight led to the design of the RGA KAN architecture. This architecture proved effective in overcoming divergence for deeper networks, with increased depth and parameter count leading to improved accuracy rather than degradation. Through the IB perspective, we linked this desirable behavior to the model’s ability to traverse all three characteristic phases of training – fitting, diffusion, and diffusion equilibrium – unlike baseline cPIKANs that tend to stall prematurely. Equipped with the proposed initialization and architecture, we then established a fixed training pipeline with adaptive components, namely RBA, RAD, causal training and learning-rate annealing. We compared RGA KANs against parameter-matched cPIKANs (also using the new initialization) and PirateNets, which are widely considered state of the art for many PDE benchmarks. Importantly, our aim was not to finely tune hyperparameters on a per-task basis but to test the performance of the proposed design in a uniform setting. Across all seven PDE benchmarks, RGA KANs outperformed both baselines, often by large margins, and remained stable in cases where the others diverged. The ablation studies further clarified the relative contribution of each adaptive method, but also revealed that in several cases the combination of the proposed initialization and architecture alone was sufficient to achieve good accuracy without divergence.
@@
Naturally, this study also has limitations. To conserve computational resources, we deliberately avoided per-task hyperparameter tuning and considered networks of medium width at 16 neurons for the final PDE benchmarks. Our focus was not to set state-of-the-art results through exhaustive tuning, but to show that the proposed initialization and architecture provide a strong and robust foundation that already outperforms both baseline cPIKANs and PirateNets under default settings. A more extensive hyperparameter search would likely further improve performance. Another limitation is that all experiments relied on first-order optimization, specifically Adam. While this is standard in the PIML literature, recent work has shown that higher-order optimizers can yield remarkable improvements [70, 71], and it would be valuable to explore the proposed architecture under such optimization regimes.
The initialization itself also opens several avenues for future work. Its basis-agnostic nature suggests that it can be applied to domains beyond PDE solving, as well as to KAN variants using other bases, which may be better suited to different tasks. Similarly, while the RGA KAN architecture was designed with PDE applications and depth-scaling in mind, its structure bears similarities to transformer architectures, hinting at its potential relevance to other application domains. Moreover, it would be interesting to test the architecture in conjunction with alternative representations of the Kolmogorov–Arnold representation theorem, such as ActNets [56] or KKANs [42], to assess whether the benefits of residual gating and adaptive training extend beyond standard KAN formulations.
Appendix A Detailed Derivations
A.1 Proposed Initialization Scheme

### v2

@@ 5.5 Advection Equation
Table 13: Results for different RGA KAN $(\alpha,\beta)$ initializations on the advection equation. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
The advection equation (see E.5 for details), especially at high transport velocities (the constant multiplying the spatial derivative of the solution field in Eq. (97)), is a challenging benchmark that typically requires specialized training strategies (e.g., learnable spatial periodical embeddings as in [PirateNets]) to obtain accurate solutions. In this work, however, we intentionally refrain from introducing any problem-specific modifications to maintain a unified training pipeline across all PDEs. To this end, we set the advection velocity to $c=20$, which remains a nontrivial setting. Table 13 shows that initializing RGA KANs with $\alpha=1$ and $\beta=1$ yields the best performance, with both the lowest mean error and the smallest variability across seeds. Among these runs, the best-performing model achieves a final relative $L^{2}$ error of $1.81\times 10^{-4}$, and its prediction is depicted in Figure 13 alongside the reference solution and absolute error.
@@
Table 14: Performance comparison on the advection equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 13. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LRA Relative $L^{2}$ Error
@@
Table 15: Ablation study on adaptive training components for the advection equation using the best $(\alpha,\beta)$ initialization from Table 13. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds.
The ablation results in Table 15 highlight the importance of adaptive training for this PDE. Training with only RBA leads to divergence, as does removing both RBA and causal training (indicated by the dash in the table), or RBA and LRA. Notably, RAD is the only adaptive strategy whose removal alongside RBA does not cause divergence, although performance is still noticeably degraded. This stands in contrast to previous benchmarks, where RAD often had the most pronounced effect, emphasizing that the relative contribution of each adaptive technique is highly problem dependent and should be evaluated individually for each PDE, as done in this study.
@@ 5.6 Helmholtz Equation
Table 16: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Helmholtz equation. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
@@
We next consider the Helmholtz equation (see E.6 for details), defined solely on a spatial domain and thus involving no initial condition. Consequently, this benchmark excludes both LRA (as the loss contains only one PDE residual term) and causal training (due to the absence of temporal dependencies). Only RBA and RAD are employed as adaptive training strategies. Table 16 shows that initializing RGA KANs with $\alpha=1$ and $\beta=1$ yields the best performance, achieving both the lowest mean relative $L^{2}$ error and the smallest variance across seeds. Notably, this configuration also attains the smallest final loss of order $\mathcal{O}(10^{-4})$, which is particularly meaningful for this benchmark: the Helmholtz equation is well known to exhibit a high and challenging loss landscape at initialization [FAIR]. The best-performing model for this configuration achieves a relative $L^{2}$ error of $1.76\times 10^{-5}$, and its prediction is shown in Figure 14.
@@
Table 17: Performance comparison on the Helmholtz equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 16. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
@@
Table 18: Ablation study on adaptive training components for the Helmholtz equation using the best $(\alpha,\beta)$ initialization from Table 16. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds.
@@ 5.7 Poisson Equation
Table 19: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Poisson equation for $\omega\in\{1,2,4\}$. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error per column is indicated in bold.
The next benchmark is the Poisson equation (see E.7 for details), which – similarly to the Helmholtz equation – is defined on a purely spatial domain and therefore involves no initial condition. As a result, only RBA and RAD are employed as adaptive training strategies. The source term is chosen such that the analytical solution is $u(x,y)=\sin(\pi\omega x)\sin(\pi\omega y)$, allowing us to systematically investigate how performance degrades as the frequency parameter $\omega$ increases. We consider $\omega\in\{1,2,4\}$ and Table 19 presents the results of the $(\alpha,\beta)$ initialization study for each $\omega$. Evidently, different initialization configurations are optimal for different frequencies, and performance deteriorates with increasing $\omega$. Moreover, the SEM for the optimal configuration grows alongside $\omega$: for $\omega=1$, it is roughly an order of magnitude lower than the mean error, whereas for $\omega=4$ it exceeds half of it, indicating increased sensitivity to initialization and optimization. For each $\omega$ value, we select the best-performing seed of the optimal configuration, achieving relative $L^{2}$ errors of $7.33\times 10^{-7}$, $2.74\times 10^{-5}$, and $3.34\times 10^{-3}$ for $\omega=1,2$, and $4$, respectively. Figure 15 shows the reference solution, prediction and absolute error for these runs.
@@
Architecture Parameters Time / Iter.
@@
Table 21: Performance comparison on the Poisson equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds for $\omega\in\{1,2,4\}$. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 19. The best performing architecture in terms of relative $L^{2}$ error per column is indicated in bold.
@@
Table 22: Ablation study on adaptive training components for the Poisson equation using the best $(\alpha,\beta)$ initialization from Table 19. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds for $\omega\in\{1,2,4\}$.
5.8 Heat Equation
For the subsequent benchmark we choose an equation with higher dimensionality, namely the (2+1)-dimensional heat equation (see E.8 for details). As indicated in Table 23, the initialization with $\alpha=0$ and $\beta=1$ yields the best performance, achieving the lowest mean relative $L^{2}$ error. The best individual run under this configuration achieved a relative $L^{2}$ error of $3.11\times 10^{-2}$. Figure 16 visualizes the model’s predictions at three temporal snapshots: $t=0$, $t=0.5$ and $t=1.0$. The absolute error field at $t=0$ reveals a localized difficulty in fitting the initial condition around $x\approx 0.85$, $y\in(0.25,0.75)$. This propagates into subsequent time instances, although it remains bounded. Despite this local artifact, the global dynamics are captured accurately, as the results demonstrate a very good overall agreement with the reference solution.
Configuration Relative $L^{2}$ Error Final Loss
$\alpha=0$, $\beta=0$ $(1.14\pm 0.26)\times 10^{-1}$ $(5.46\pm 2.33)\times 10^{-5}$
$\alpha=1$, $\beta=0$ $(4.91\pm 0.77)\times 10^{-2}$ $(2.71\pm 0.31)\times 10^{-5}$
$\mathbf{\boldsymbol{\alpha}=0}$, $\mathbf{\boldsymbol{\beta}=1}$ $\mathbf{(3.73\pm 0.36)\times 10^{-2}}$ $\mathbf{(2.57\pm 0.32)\times 10^{-5}}$
$\alpha=1$, $\beta=1$ $(6.44\pm 1.39)\times 10^{-2}$ $(3.78\pm 2.93)\times 10^{-5}$
Table 23: Results for different RGA KAN $(\alpha,\beta)$ initializations on the heat equation. Reported values are mean $\pm$ SEM over three seeds. The best performing configuration in terms of relative $L^{2}$ error is indicated in bold.
Figure 16: Reference solution (left column), RGA KAN prediction (middle column) and absolute error (right column) for the heat equation, shown for the random seed corresponding to the best-performing model instance at three different snapshots: $t=0$ (top row), $t=0.5$ (middle row), $t=1.0$ (bottom row).
The comparative results across architectures are presented in Table 24. Consistent with previous benchmarks, the cPIKAN model struggles to learn the solution, exhibiting a high mean relative error and significant variance. PirateNet offers a marked improvement over cPIKAN but fails to match the precision of the proposed method. RGA KANs achieve the lowest error and demonstrate superior consistency, with a SEM that is substantially lower than that of the baseline architectures.
Architecture Parameters Relative $L^{2}$ Error Time / Iter.
cPIKAN 18,397 $(5.12\pm 2.52)\times 10^{-1}$ 3.60 ms
PirateNet 19,246 $(1.78\pm 0.82)\times 10^{-1}$ 3.78 ms
RGA KAN 18,502 $\mathbf{(3.73\pm 0.36)\times 10^{-2}}$ 5.52 ms
Table 24: Performance comparison on the heat equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses the best $(\alpha,\beta)$ initialization from Table 23. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
RBA RAD Causal LRA Relative $L^{2}$ Error
$(3.73\pm 0.36)\times 10^{-2}$
$(3.23\pm 0.94)\times 10^{-2}$
$(7.92\pm 1.01)\times 10^{-2}$
$(8.68\pm 1.65)\times 10^{-2}$
$(1.13\pm 0.11)\times 10^{-1}$
$(1.05\pm 0.18)\times 10^{-1}$
Table 25: Ablation study on adaptive training components for the heat equation using the best $(\alpha,\beta)$ initialization from Table 23. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error are mean $\pm$ SEM over three seeds.
Finally, Table 25 summarizes the ablation study for this benchmark. A notable observation here is the critical role of RBA: the configuration using only RBA performs competitively with the fully adaptive setup, yielding a slightly lower mean error but with higher variance. Conversely, disabling RBA roughly doubles the error. Further removing either the causal training or LRA mechanisms leads to a significant degradation in performance, increasing the error by an order of magnitude.
5.9 Navier-Stokes Equation
The final benchmark is the Navier-Stokes equation (see E.9 for details), modeling viscous fluid dynamics on a torus. Following established practices in the literature [jaxpi], we introduce a weighting factor of $100$ for the continuity equation residual and initialize the initial condition weight $\lambda_{\text{ic}}$ to $10^{5}$ to prioritize the fitting of the initial snapshot. We adopt these settings as the default for our experiments to ensure a fair comparison with baselines, though we explicitly investigate their necessity in our ablation study.
Configuration Relative $L^{2}$ Error Final Loss
$\alpha=0$, $\beta=0$ $u$: $(1.08\pm 0.40)\times 10^{-2}$ $(5.14\pm 0.20)\times 10^{-2}$
$v$: $(7.01\pm 2.26)\times 10^{-3}$
$w$: $(1.54\pm 0.15)\times 10^{-3}$
$\alpha=1$, $\beta=0$ $u$: $\mathbf{(8.90\pm 3.44)\times 10^{-3}}$ $(6.18\pm 0.16)\times 10^{-2}$
$v$: $\mathbf{(4.78\pm 1.16)\times 10^{-3}}$
$w$: $(1.86\pm 0.25)\times 10^{-3}$
$\alpha=0$, $\beta=1$ $u$: $(1.33\pm 0.50)\times 10^{-2}$ $(2.81\pm 0.08)\times 10^{-2}$
$v$: $(8.37\pm 3.04)\times 10^{-3}$
$w$: $\mathbf{(1.60\pm 0.23)\times 10^{-3}}$
$\alpha=1$, $\beta=1$ $u$: $(1.51\pm 0.51)\times 10^{-2}$ $(2.81\pm 0.08)\times 10^{-2}$
$v$: $(8.11\pm 2.49)\times 10^{-3}$
$w$: $(7.73\pm 2.52)\times 10^{-3}$
Table 26: Results for different RGA KAN $(\alpha,\beta)$ initializations on the Navier-Stokes equation. Reported values are mean $\pm$ SEM over three seeds. Errors are reported separately for velocity components ($u,v$) and vorticity ($w$). The best performance per field is indicated in bold.
We first examine the impact of initialization on model performance. While the vorticity field ($w$) is the standard metric reported for this benchmark, we also provide errors for the velocity components ($u,v$) in Table 26 for completeness. The results show that while the configuration $\alpha=0,\beta=1$ yields the lowest error for vorticity, the configuration $\alpha=1,\beta=0$ achieves significantly better accuracy for the velocity fields $u$ and $v$ with only a marginal trade-off in vorticity performance. Consequently, we select $\alpha=1,\beta=0$ as the optimal configuration for the subsequent comparisons. Figure 17 visualizes the predictions for this model at $t=0.1$, for a run achieving relative $L^{2}$ errors of $6.02\times 10^{-3}$ for $u$, $4.93\times 10^{-3}$ for $v$, and $1.63\times 10^{-3}$ for $w$.
Figure 17: Reference solution (left column), RGA KAN prediction (middle column) and absolute error (right column) for the solutions to the Navier-Stokes equation, shown for the random seed corresponding to the best-performing model instance at the final temporal snapshot $t=0.1$. The top, middle and bottom row corresponds to $u$, $v$ and $w$, respectively.
Table 27 compares the RGA KAN against the cPIKAN and PirateNet baselines. A notable observation here is the shift in computational cost compared to previous benchmarks. While cPIKAN remains the fastest, the training times for PirateNet and RGA KAN are now comparable. This is due to the heavy computational load of the Navier-Stokes loss function, which involves multiple high-order derivatives. In this regime, the gating mechanisms shared by both PirateNet and RGA KAN become the primary bottleneck, rendering the difference between computing KAN basis functions and MLP operations less significant. Despite the similar cost, the performance gap is substantial: RGA KAN outperforms PirateNet by an order of magnitude on the velocity fields and achieves a vorticity error nearly three times lower.
Architecture Params Relative $L^{2}$ Error Time / Iter.
cPIKAN 18,578 $u$: $(3.05\pm 0.40)\times 10^{-2}$ 15.88 ms
$v$: $(2.87\pm 0.32)\times 10^{-2}$
$w$: $(9.25\pm 0.66)\times 10^{-3}$
PirateNet 19,300 $u$: $(5.00\pm 1.22)\times 10^{-2}$ 24.85 ms
$v$: $(2.31\pm 0.12)\times 10^{-2}$
$w$: $(4.19\pm 0.73)\times 10^{-3}$
RGA KAN 18,662 $u$: $\mathbf{(8.90\pm 3.44)\times 10^{-3}}$ 26.40 ms
$v$: $\mathbf{(4.78\pm 1.16)\times 10^{-3}}$
$w$: $\mathbf{(1.86\pm 0.25)\times 10^{-3}}$
Table 27: Performance comparison on the Navier-Stokes equation across different architectures. Reported values are mean $\pm$ SEM over three random seeds. The RGA KAN row uses $(\alpha,\beta)=(1.0,0.0)$. The best performing architecture in terms of relative $L^{2}$ error is indicated in bold.
Weights RBA RAD Causal LRA Relative $L^{2}$ Error
$(1.86\pm 0.25)\times 10^{-3}$
$(2.96\pm 0.18)\times 10^{-3}$
$(1.48\pm 0.19)\times 10^{-2}$
$(1.58\pm 0.08)\times 10^{-3}$
$(7.69\pm 0.75)\times 10^{-3}$
$(1.64\pm 0.11)\times 10^{-3}$
$(6.57\pm 0.27)\times 10^{-2}$
Table 28: Ablation study on adaptive training components for the Navier-Stokes equation using $(\alpha,\beta)=(1.0,0.0)$. Each row corresponds to a different combination of enabled (✓) or disabled (✗) components. Reported values for relative $L^{2}$ error of the vorticity field ($w$) are mean $\pm$ SEM over three seeds.
Finally, we extend the ablation study in Table 28 to investigate the role of the manual loss weights. We find that removing these heuristic weights (“Weights” column) increases the error, confirming their utility; however, the model does not diverge. Crucially, even without manual weighting, the RGA KAN yields a lower error than the weighted cPIKAN and PirateNet models. Regarding the adaptive components, we observe a distinct behavior compared to, e.g., the heat equation: here, RBA appears to hinder fine-scale convergence. Removing RBA leads to the lowest overall error, whereas relying on RBA alone (row 3) degrades performance by an order of magnitude.
@@ 6 Conclusion and Outlook
In this work, we studied the deep training of Chebyshev-based KANs with the goal of improving their stability and accuracy in PDE benchmarks under a uniform training setup. We began by examining their initialization properties and proposed a Glorot-like initialization scheme that is basis-agnostic (as demonstrated by its application to sine-based KANs within the RGA KAN architecture). Preliminary results on function fitting and PDE tasks showed that this initialization alone significantly improved training outcomes compared to the default initialization, in some cases by several orders of magnitude. For certain benchmarks such as Burgers’ equation, the proposed initialization was sufficient to train deeper models successfully, whereas for others, namely Allen–Cahn, it fell short. This observation made it clear that initialization alone was not enough to fully address the depth-scaling issue.
@@
Motivated by this, and inspired by PirateNets, we analyzed the properties of Chebyshev KANs at initialization and observed strong parallels to standard MLPs. This insight led to the design of the RGA KAN architecture. This architecture proved effective in overcoming divergence for deeper networks, with increased depth and parameter count leading to improved accuracy rather than degradation. Through the IB perspective, we linked this desirable behavior to the model’s ability to traverse all three characteristic phases of training – fitting, diffusion, and diffusion equilibrium – unlike baseline cPIKANs that tend to stall prematurely. Equipped with the proposed initialization and architecture, we then established a fixed training pipeline with adaptive components and compared RGA KANs against parameter-matched cPIKANs (also using the new initialization) and PirateNets, which are widely considered state of the art for many PDE benchmarks. Importantly, our aim was not to finely tune hyperparameters on a per-task basis but to test the performance of the proposed design in a uniform setting. Across all PDE benchmarks, RGA KANs outperformed both baselines, often by large margins, and remained stable in cases where the others diverged. The ablation studies further clarified the relative contribution of each adaptive method, but also revealed that in several cases the combination of the proposed initialization and architecture alone was sufficient to achieve good accuracy without divergence.
Regarding the architecture’s specific configuration, we extensively tested four distinct initialization pairs for the gating parameters $\alpha$ and $\beta$, namely combinations of 0 and 1. While this sweep highlights how the network responds to different initial effective depths, it does not imply that a parametric search is strictly necessary for every new problem. In fact, our experiments suggest that initializing with mixed states – either $(\alpha,\beta)=(1,0)$ or $(0,1)$ – consistently yields superior performance compared to baseline architectures. The choice of 0 and 1 as initialization values carries physical significance: $(0,0)$ forces the network to adaptively grow its effective depth from a shallow state, while $(1,1)$ initializes the network at full depth, relying on the robustness of the proposed initialization scheme. It is important to clarify that while we initialize these parameters to discrete extremes, they are defined as continuous, trainable variables. During optimization, they naturally evolve into floating-point values.
While performance and stability are paramount, computational complexity remains a critical factor. Under fair parameter-matched settings, the time complexity of the RGA KAN is generally higher than that of PirateNets or cPIKANs. As observed in simpler benchmarks, this overhead stems from the simultaneous computation of multiple basis functions combined with the additional operations within the gating mechanisms. However, this dynamic changes as the problem complexity increases. As demonstrated in the Navier-Stokes benchmark, where the loss function involves multiple high-order spatial derivatives, the primary computational bottleneck shifts toward the backpropagation through the dense gating mechanisms. This explains why the training times of RGA KANs and PirateNets (both of which utilize gating) converge to similar values in this regime.
@@
Naturally, this study also has limitations. To conserve computational resources, we deliberately avoided per-task hyperparameter tuning and considered networks of medium width at 16 neurons for the final PDE benchmarks. Our focus was not to set state-of-the-art results through exhaustive tuning, but to show that the proposed initialization and architecture provide a strong and robust foundation that already outperforms both baseline cPIKANs and PirateNets under default settings. A more extensive hyperparameter search would likely further improve performance. Another limitation is that all experiments relied on first-order optimization, specifically Adam. While this is standard in the PIML literature, recent work has shown that higher-order optimizers can yield remarkable improvements [SSBroyden, SOAP], and it would be valuable to explore the proposed architecture under such optimization regimes.
The proposed framework opens several avenues for future work. The initialization’s basis-agnostic nature suggests potential applications to KAN variants using other bases, while the RGA KAN’s transformer-like structure hints at relevance in broader domains. A particularly promising direction is the extension of this architecture to operator learning [DeepONet]. While this paradigm focuses on learning mappings between function spaces rather than solving single instances, it shares the fundamental challenge of approximating PDE solutions. Given that KAN-based architectures have already shown promise in this field [DeepOKAN, KANO, KKANs], utilizing the proposed architecture in that context has potential. Finally, it would be interesting to test the architecture using KANs that are based on alternative representations of the Kolmogorov–Arnold theorem, such as ActNets [ActNet] or KKANs [KKANs], to assess whether the benefits of residual gating and adaptive training extend beyond the formulation used in this work.
Appendix A Adaptive Training Methods
A.1 Collocation Points Resampling
The selection of collocation points used to enforce the PDE can be performed in a one-time manner, where a fixed set of training points is used throughout the entire training process, either with or without mini-batching. However, periodically resampling the collocation points during training can serve as a regularization technique, improving the network’s ability to generalize to spatiotemporal regions that were not explicitly sampled. Moreover, if the resampling process is not purely random but instead adaptive – e.g., guided by residuals or other heuristics – it has been shown to significantly enhance the final accuracy of the trained network [COLLOCS1, COLLOCS2, COLLOCS3].
For this study, we adopt the residual-based adaptive distribution (RAD) technique [RAD]. In particular, an initial dense pool of $N_{\text{pool}}$ collocation points, $\left\{\left(t_{\text{pool}}^{i},\mathbf{x}_{\text{pool}}^{i}\right)\right\}_{i=1}^{N_{\text{pool}}}$, is generated from a uniform grid, and training is performed on a dynamically selected subset of $N_{\text{pde}}\ll N_{\text{pool}}$ points. These points are periodically resampled from the pool according to the probability density function
A.2 Global Loss Weighting
Among the numerous proposed ways to address this issue [WangNTK, KKANs, AAHWeights], we choose the learning-rate annealing (LRA) algorithm introduced in [Wang1]. During training, the weight adjustment is guided by the computation
[…]
A.3 Causal Training
A common challenge in training neural networks to solve time-dependent PDEs is the violation of causality. Since collocation points are sampled from the entire temporal domain, the network may unintentionally minimize residuals associated with future states before adequately minimizing those corresponding to past states. To mitigate this issue, following [Causal], we partition the temporal domain into $M$ sequential segments of equal length and introduce temporal weights $\left\{w_{i}\right\}_{i=1}^{M}$, modifying Eq. (7) as
A.4 Residual-Based Attention
The LRA algorithm introduced in Section A.2 is essentially a global loss weighting scheme, while causal training corresponds to a batch-wise weighting strategy, where groups of collocation points share the same weights. However, point-wise multipliers have also demonstrated significant success in PIML [SA] and have been integral to works achieving state-of-the-art results [KKANs, BRDR]. In such methods, the individual loss terms in Eqs. (7)-(9) are modified as
For this study, we adopt the residual-based attention (RBA) method introduced in [RBA], where all local weights are initially set to 1 and updated at each training iteration according to
where $\gamma\geq 0$, $\eta\geq 0$ are hyperparameters of the method.
In theory, while RBA is not inherently incompatible with causal training – as the batch-wise weights can be computed using the modified loss of Eq. (55) – the same does not hold for RAD as presented in A.1. Since RAD involves resampling collocation points, it removes the one-to-one correspondence between them and their local weights. In order to apply these methods in conjunction, we adopt the following strategy: instead of assigning local weights solely to the currently selected $N_{\text{pde}}$ training points, we define the RBA weights over the entire pool, i.e., $\alpha_{\text{pool}}^{i}=1$ for $i=1,\dots,N_{\text{pool}}$ at initialization. These weights are updated at each training iteration according to Eq. (56), using the residuals evaluated at the corresponding points. Then, during each application of RAD, the probability density function in Eq. (50) is modified by incorporating the RBA weights as multiplicative factors on the residuals:
Appendix B Detailed Derivations
B.1 Proposed Initialization Scheme